OpenStudy (anonymous):

What is the equation of a line tangent to y= sin x + cos x at ( pi/2,1) ?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

is the gradient

OpenStudy (ksaimouli):

what math r uin

OpenStudy (anonymous):

Calculus

OpenStudy (ksaimouli):

bc or ab which grade r u in plz iam curious to know

OpenStudy (anonymous):

BC junior

OpenStudy (ksaimouli):

wow ru in collage

OpenStudy (anonymous):

not yet

OpenStudy (ksaimouli):

high school

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

\[y \prime = cosx - sinx ,,, => \cos(\pi/2) = 0 , sinx (\pi/2) = 1\] so the answer is -1

OpenStudy (anonymous):

thank you